# Smallest Suffixient Sets as a Repetitiveness Measure

### v2

@@ Abstract
A suffixient set is a novel combinatorial object that captures the essential information of repetitive strings in a way that, provided with a random access mechanism, supports various forms of pattern matching. In this paper, we study the size $\chi$ of the smallest suffixient set as a repetitiveness measure: we place it between known measures and study its sensitivity to various string operations. As a corollary of our results, we give a simple online algorithm to compute smallest suffixient sets.
@@ 1 Introduction
Our final contribution is a new, extremely simple, online algorithm to compute smallest suffixient sets (and thus $\chi$). While there already exist efficient algorithms to do this [5], our new algorithm processes the string left to right and, at any point, can exhibit a smallest suffixient set for the string it has just consumed. Just as online suffix tree construction [33], our algorithm uses $O(n)$ space and worst-case time in the transdichotomous RAM model over polynomial-size integer alphabets. The best previous algorithm obtains the same result [5], but it is not online and starts from the suffix array and other components of the suffix tree, whereas ours starts from the text and builds the suffix tree at the same time. All linear-time suffix tree construction algorithms run under the same model of computation.
@@ 2 Preliminaries
A string $w[1\mathinner{.\,.}n]$ (or simply $w$ if it is clear from the context) of length $|w|=n$ over the alphabet $\Sigma$ is a sequence $w[1]w[2]\cdots w[n]$ of symbols where $w[i]\in\Sigma$ for all $i\in[1,n]$. The empty string of length 0 is denoted $\epsilon$. We denote by $\Sigma^{*}$ the set of all strings over $\Sigma$. Additionally, we let $\Sigma^{+}=\Sigma^{*}\setminus\{\epsilon\}$ and $\Sigma^{k}=\{w\in\Sigma^{*}\,|\,|w|=k\}$. We denote by $w[i\mathinner{.\,.}j]$ the substring $w[i]w[i+1]\cdots w[j]$. If $x=x[1\mathinner{.\,.}n]$ and $y=y[1\mathinner{.\,.}m]$ are strings, we define the concatenation operation applied on $x$ and $y$, as the string obtained by juxtaposing these two strings, that is, $x\cdot y=x[1]x[2]\cdots x[n]y[1]\cdots y[m]=xy$. A string $x$ is a substring of $w$ if $w=yxz$ for some $y,z\in\Sigma^{*}$. A string $x$ is a prefix of $w$ if $w=xy$ for some $y\in\Sigma^{*}$. Analogously, $x$ is a suffix of $w$ if $w=yx$ for some $y\in\Sigma^{*}$. We say that substrings, prefixes, and suffixes are non-trivial if they are different from $w$ and $\epsilon$. The set of substrings of $w$ is denoted by $\mathcal{F}_{w}$. We also let $\mathcal{F}_{w}(k)=\mathcal{F}_{w}\cap\Sigma^{k}$. The reverse of a finite string $w$ is the string $w^{R}=w[n]\cdot w[n-1]\cdots w[1]$. We denote by $\mathcal{R}(w)$ the multiset of rotations of $w[1\mathinner{.\,.}n]$, that is, $\mathcal{R}(w)=\{w[i+1\mathinner{.\,.}n]w[1\mathinner{.\,.}i]\,|\,i\in[1\mathinner{.\,.}n]\}$. The Burrows-Wheeler transform (BWT) of a string $w$, denoted $\mathtt{BWT}(w)$, is the transformation of $w$ obtained by collecting the last symbol of all rotations in $\mathcal{R}(w)$ in lexicographic order. The BWT matrix $B(w)$ of $w$ is the $(n\times n)$-matrix where the $i$-th row is the $i$-th rotation of $w$ in lexicographic order.
@@ 4 Sensitivity of $\chi$ to String Operations
The sensitivity to string operations has been studied for many repetitiveness measures [1, 9, 10, 14, 15, 24, 29, 30]. It is desirable for a repetitiveness measure to not change much upon small changes in the sequence. Some repetitiveness measures are resistant to edit operations. For instance, $b$, $z$ and $g$ can only increase by a multiplicative constant after an edit operation [1], though they can increase only by a $O(1)$ additive factor when prepending or appending a character. On the other hand, $r$ can increase by a $\Theta(\log n)$ factor when appending a character [15, Prop. 37]. Other results have been obtained concerning more complex string operations, like reversing a string [14], or applying a string morphism [9, 10].
In this section we study how $\mathtt{sre}$ and $\chi$ behave in this respect. We start by proving the following useful lemma.
@@
Lemma 2 yields a remarkably simple algorithm to compute smallest suffixient sets, which we detail in Section 6.
@@
The lower bound follows from Lemma 1. For the upper bound, let cxacxa be the smallest prefix of $cw$ that is not a right-extension of $w$, but is a right-extension of $cw$ (if it exists). This means that cxacxa does not appear in $w$ (otherwise, it would be a right-extension of $w)$, so no prefix of $cw$ of length $|cxa|$ or more is right-maximal. Hence, all prefixes of $cw$ shorter than cxacxa were already right-extensions, and all prefixes longer than cxacxa cannot be right-extensions. Therefore, cxacxa together with some cxbcxb appearing in $w$, are the only possible new right-extensions in $cw$ with respect to ww.∎
By letting $c=\text{\tt\textdollar}\not\in\mathcal{F}_{w}$ in Lemma 2, we relate $\chi$ to $\mathtt{sre}$ (note that $\chi$ is always at least $\mathtt{sre}+1$ because of the new super-maximal extension ending with \$). This makes clear the relation between Combinatorics on words [21] with suffixient sets, via the common notion of right-special factors (what we call here right-maximal substrings).
@@
Now we study how much $\mathtt{sre}(w)$ can vary upon edit operations in arbitrary positions, rotations, and reversals. We will use the following famous string family.
@@
A binary de Bruijn sequence of order $k>0$ [2] contains every binary string in $\{\texttt{a},\texttt{b}\}^{k}$ as a substring exactly once. The length of these strings is $n=2^{k}+(k-1)$. The set of binary de Bruijn sequences of order $k$ is $\mathtt{dB}(k)$.
@@
This sums to a total of $4k$ super-maximal extensions in $w_{k}^{R}$. Thus, $\mathtt{sre}(w_{k})-\mathtt{sre}(w_{k}^{R})=(5k-1)-4k=k-1$. ∎
We give an example of the words $w_{k}$ and $w_{k}^{R}$ of Lemma 7, and their super-maximal right-extensions.
@@
Formally, the additive sensitivity of a measure of repetitiveness $\mu$ to a string operation $\rho$ can be defined as a function $AS_{\mu,\rho}:\mathbb{Z}^{+}\rightarrow\mathbb{R}$, where $AS_{\mu,\rho}(n)=\max_{w\in\Sigma^{n}}(\max_{w^{\prime}\in\rho(w)}(\mu(w^{\prime}))-\mu(w))$, that is the maximum achievable difference among all the strings. Overall, we obtain the following result on the additive sensitivity of $\mathtt{sre}$, which, by Corollary 1, can be written in terms of $\chi$.
@@
Finally, we show upper bounds on the sensitivity of $\chi$ to string operations.
@@
$\chi(w^{\prime})-\chi(w)=O\left(\delta\max\left(1,\log(n/\delta\log\delta)\right)\log\delta\right)\text{ and }$
$\chi(w^{\prime})\hskip 3.0pt/\hskip 4.0pt\chi(w)=O\left(\max\left(1,\log(n/\delta\log\delta)\right)\log\delta\right).$
@@
To prove our thesis, we rely on the relations $\delta\leq\chi\leq 2\overline{r}$ [4] and $r=O(\delta\max(1,\log(n/\delta\log\delta))\log\delta)$ [17]. Moreover, since the multiplicative sensitivity of the measure $\delta$ to any of the string operations is $O(1)$ [1], for any $w\in\Sigma^{*}$ it holds $\overline{r}(w)=r(w^{R})=O(\delta\max(1,\log(n/\delta\log\delta))\log\delta)$. The thesis follows by considering the worst case, that is $\chi(w)=\Theta(\delta)$ and $\chi(w^{\prime})=\Theta(\delta\max(1,\log(n/\delta\log\delta))\log\delta)$.∎
@@ 5.1 Proving $\chi=O(r)$
Let $x_{i}$ denotes the $i$th rotation of $w\text{\tt\textdollar}$ in lexicographic order, for each $i\in[1\mathinner{.\,.}|w|+1]$, and let $u_{i}$ be the longest common prefix between the rotations $x_{i},x_{i+1}$, for each $i\in[1\mathinner{.\,.}|w|]$. We further define $s:[1\mathinner{.\,.}n+1]\rightarrow[0\mathinner{.\,.}n]$ as $s(i)=j$ if $x_{i}=w[j+1\mathinner{.\,.}|w|]\text{\tt\textdollar}w[1\mathinner{.\,.}j]$, i.e., the number of cyclic shift to the right required to transform $x_{i}$ into $w\text{\tt\textdollar}$. [^2] The function $s$ mimics the well-known Suffix Array [23], here omitted for simplicity of exposition. As the symbol \$ occurs only once in $w\text{\tt\textdollar}$, the function $s$ is bijective.
@@
that is the set of positions where the occurrences of the right-extensions $u_{i}c_{1}$ and $u_{i}c_{2}$ end in $w\text{\tt\textdollar}$, where $u_{i}c_{1}$ and $u_{i}c_{2}$ are the prefix of $x_{i}$ and $x_{i+1}$ respectively, for some $c_{1},c_{2}\in\Sigma$ such that $c_{1}<c_{2}$. It follows by construction that the set $S$ is a suffixient set of $w\text{\tt\textdollar}$.
@@ 5.2 A Family with $\chi=o(v)$ (and thus $o(r)$)
An infinite string w is episturmian if it has at most one right-maximal substring of each length and its set of substrings is closed under reversal, that is, $\mathcal{F}_{\textbf{w}}=\mathcal{F}_{\textbf{w}}^{R}$. It is standard episturmian (or epistandard) if, in addition, all the right-maximal substrings of w are of the form $\textbf{w}[1\mathinner{.\,.}i]^{R}$ with $i\geq 0$, i.e., they are the reverse of some prefix of w.
@@
Let $\textbf{w}\in\Sigma^{\omega}$ be an episturmian word with $\sigma\geq 2$. Then, $\mathtt{sre}(\textbf{w}[i\mathinner{.\,.}j])\leq\sigma$ for $i,j\geq 0$.
@@
Let w be an epistandard word. The right-extensions $x_{1},x_{2},\dots$ ending with $a\in\Sigma$ form a suffix-chain where each $x_{i}$ is a suffix of $x_{i+1}$. There is one of those suffix-chains for each character $a\in\Sigma$.
@@
When considering substrings of w, the super-maximal right-extension in $\textbf{w}[i\mathinner{.\,.}j]$ ending with $a\in\Sigma$ is the longest right-extension of w ending with $a$ that remains a right-extension in $\textbf{w}[i\mathinner{.\,.}j]$. It follows that for any substring $\textbf{w}[i\mathinner{.\,.}j]$ of any episturmian word w, $\mathtt{sre}\leq\sigma$. ∎
@@ 6 Online Computation of Smallest Suffixient Sets
As an application of Lemma 2, we show that Ukkonen’s linear-time online construction of suffix trees [33] can be easily modified to compute smallest suffixient sets within the same space and time complexity.
The suffix tree of a text $T$ is a tree of size $O(|T|)$ where edges are labeled by nonempty substrings of $T$ (the label $T[i\mathinner{.\,.}j]$ is indicated by the pair $(i,j)$). Every node has at least two children, and the labels of edges to any two children must differ in their first symbol. If node $x$ has a child node $y$ by an edge labeled $(i,j)$, we say that $y$ is a child by label $T[i]$. Each node $x$ is said to be the locus of the string obtained by concatenating the labels of the edges that lead from the root to $x$. The string depth of a node is the length of its locus. The tree leaves are the loci of suffixes $T[i\mathinner{.\,.}]$, whereas internal nodes are loci of strings that occur at least twice in $T$. Edges to leaves have labels of the form $(i,\infty)$, which means the second component is always the last position processed of $T$. Suffix tree nodes also have so-called suffix links, which lead from a node that is the loci of a string $cu$, for $c\in\Sigma$, to the loci of string $u$ (which always exists: if $cu$ occurs more than once, so does $u$). Finally, we extend the concept of suffix tree nodes to implicit nodes, which are virtual nodes with one child, assumed to exist along edges: if $y$ is the child of $x$ by an edge labeled $(i,j)$ and $x$ is the locus of string $u$, then $(x,(i,p))$, for $i\leq p<j$, is an implicit nodes whose locus is $u\cdot T[i\mathinner{.\,.}p]$. Knowing their represented string, suffix links are also defined from implicit nodes (those can lead to explicit or to implicit nodes).
Ukkonen’s algorithm builds the suffix tree of $T$ such that, after having processed any prefix $w$ of $T$, it has built the suffix tree of $w$. To prevent special cases, it assumes that the root is in fact a child of a special node $\bot$, with edges to the root labeled $c$ for every $c\in\Sigma$; the suffix link of the root is $\bot$. We do not aim at a full explanation of the algorithm, but just highlight some of its relevant properties. Let the prefix $w$ of $T$ be followed by $c\in\Sigma$. The algorithm maintains pointers to two (possibly implicit) nodes of the suffix tree of $w$:

### v3

@@ Abstract
A suffixient set is a novel combinatorial object that captures the essential information of repetitive strings in a way that, provided with a random access mechanism, supports various forms of pattern matching. In this paper, we study the size $\chi$ of the smallest suffixient set as a repetitiveness measure: we place it between known measures and study its sensitivity to various string operations. As a corollary of our results, we give simple online algorithms to compute smallest suffixient sets.
@@ 1 Introduction
Our final contribution are new, extremely simple, online algorithms to compute smallest suffixient sets (and thus $\chi$), scanning the text left to right or right to left. While there already exist efficient algorithms to do this [5], our new algorithm can, at any point, exhibit a smallest suffixient set for the string it has just consumed. Just as online suffix tree constructions [34, 33], our algorithm uses $O(n)$ space and worst-case time in the transdichotomous RAM model over polynomial-size integer alphabets. The best previous algorithm obtains the same result [5], but it is not online and starts from the suffix array and other components of the suffix tree, whereas ours starts from the text and builds the suffix tree at the same time. All linear-time suffix tree construction algorithms run under the same model of computation.
@@ 2 Preliminaries
A string $w[1\mathinner{.\,.}n]$ (or simply $w$ if it is clear from the context) of length $|w|=n$ over the alphabet $\Sigma$ is a sequence $w[1]w[2]\cdots w[n]$ of symbols where $w[i]\in\Sigma$ for all $i\in[1,n]$. The empty string of length 0 is denoted $\epsilon$. We denote by $\Sigma^{*}$ the set of all strings over $\Sigma$. Additionally, we let $\Sigma^{+}=\Sigma^{*}\setminus\{\epsilon\}$ and $\Sigma^{k}=\{w\in\Sigma^{*}\mid|w|=k\}$. We denote by $w[i\mathinner{.\,.}j]$ the substring $w[i]w[i+1]\cdots w[j]$. If $x=x[1\mathinner{.\,.}n]$ and $y=y[1\mathinner{.\,.}m]$ are strings, we define the concatenation operation applied on $x$ and $y$, as the string obtained by juxtaposing these two strings, that is, $x\cdot y=x[1]x[2]\cdots x[n]y[1]\cdots y[m]=xy$. A string $x$ is a substring of $w$ if $w=yxz$ for some $y,z\in\Sigma^{*}$. A string $x$ is a prefix of $w$ if $w=xy$ for some $y\in\Sigma^{*}$. Analogously, $x$ is a suffix of $w$ if $w=yx$ for some $y\in\Sigma^{*}$. We say that substrings, prefixes, and suffixes are non-trivial if they are different from $w$ and $\epsilon$. The set of substrings of $w$ is denoted by $\mathcal{F}_{w}$. We also let $\mathcal{F}_{w}(k)=\mathcal{F}_{w}\cap\Sigma^{k}$. The reverse of a finite string $w$ is the string $w^{R}=w[n]\cdot w[n-1]\cdots w[1]$. We denote by $\mathcal{R}(w)$ the multiset of rotations of $w[1\mathinner{.\,.}n]$, that is, $\mathcal{R}(w)=\{w[i+1\mathinner{.\,.}n]w[1\mathinner{.\,.}i]\mid i\in[1\mathinner{.\,.}n]\}$. The Burrows-Wheeler transform (BWT) of a string $w$, denoted $\mathtt{BWT}(w)$, is the transformation of $w$ obtained by collecting the last symbol of all rotations in $\mathcal{R}(w)$ in lexicographic order. The BWT matrix $B(w)$ of $w$ is the $(n\times n)$-matrix where the $i$-th row is the $i$-th rotation of $w$ in lexicographic order.
@@ 4 Sensitivity of $\chi$ to String Operations
The sensitivity to string operations has been studied for many repetitiveness measures [1, 9, 10, 14, 15, 24, 29, 30]. A robust repetitiveness measure should not change much upon small changes in the sequence. For instance, $b$, $z$, and $g$ can increase only by a multiplicative constant after an edit operation [1], and they can increase only by a constant additive factor when prepending or appending a character. On the other hand, $r$ can increase by a $\Theta(\log n)$ factor when appending a character [15, Prop. 37]. Other results have been obtained concerning more complex string operations, like reversing a string [14], or applying a string morphism [9, 10].
In this section we study how $\mathtt{sre}$ and $\chi$ behave in this respect.
4.1 Appending and prepending symbols
We first show that $\mathtt{sre}(w)$ and $\chi(w)$ increase or decrease by only additive constants when adding or removing symbols at the extremes of $w$. We start by proving the following useful lemma.
@@ 4.1 Appending and prepending symbols
The lower bound follows from Lemma 1. For the upper bound, let cxacxa be the smallest prefix of $cw$ that is not a right-extension of $w$, but is a right-extension of $cw$ (if it exists). This implies that there exists $b\neq a$ such that $cxb\in\mathcal{F}_{w}$ and $cxa\not\in\mathcal{F}_{w}$ (otherwise, cxacxa would be a right-extension of $w)$, so no prefix of $cw$ of length $|cxa|$ or more is right-maximal, and thus no prefix longer than cxacxa can be a right-extension. By the minimality of cxacxa, all prefixes of $cw$ shorter than cxacxa either are not right-extensions of $cw$, or are right-extensions of both $w$ and $cw$. Therefore, cxacxa together with some cxbcxb appearing in $w$, are the only possible new right-extensions in $cw$ with respect to ww.∎
Lemmas 2 and 3 yield a remarkably simple algorithms to compute smallest suffixient sets, which we detail in Section 6.
By letting $c=\text{\tt\textdollar}\not\in\mathcal{F}_{w}$ in Lemma 2, we relate $\chi$ to $\mathtt{sre}$ (note that $\chi$ is always at least $\mathtt{sre}+1$ because of the new super-maximal extension ending with \$). This makes clear the relation between Combinatorics on words [21] with suffixient sets, via the common notion of right-special factors (which we call here right-maximal substrings).
@@
4.2 Arbitrary edit operations and rotations
Now we study how much $\mathtt{sre}(w)$ can vary upon edit operations at arbitrary positions and rotations. We will use the following famous string family.
@@ 4.2 Arbitrary edit operations and rotations
A binary de Bruijn sequence of order $k>0$ [2] contains every binary string in $\{\texttt{a},\texttt{b}\}^{k}$ as a substring exactly once. The length of this sequence is $n=2^{k}+(k-1)$. The set of binary de Bruijn sequences of order $k$ is $\mathtt{dB}(k)$.
@@ 4.3 Reversals
This sums to a total of $4k$ super-maximal extensions in $w_{k}^{R}$. Thus, $\mathtt{sre}(w_{k})-\mathtt{sre}(w_{k}^{R})=(5k-1)-4k=k-1$. This is $\Theta(\sqrt{|w_{k}|})$ since $|w_{k}|=2k^{2}+4k$. ∎
As the construction is quite complex, we give an example of the words $w_{k}$ and $w_{k}^{R}$ of Lemma 7, and their super-maximal right-extensions.
@@
4.4 Lower and upper bounds on the sensitivity of $\chi$
Formally, the additive sensitivity of a measure of repetitiveness $\mu$ to a string operation $\rho$ can be defined as a function $AS_{\mu,\rho}:\mathbb{Z}^{+}\rightarrow\mathbb{R}$, where $AS_{\mu,\rho}(n)=\max_{w\in\Sigma^{n}}(\max_{w^{\prime}\in\rho(w)}(\mu(w^{\prime}))-\mu(w))$, that is, the maximum achievable difference among all the strings. Overall, we obtain the following result on the additive sensitivity of $\mathtt{sre}$, which, by Corollary 1, can be written in terms of $\chi$.
@@ 4.4 Lower and upper bounds on the sensitivity of $\chi$
Finally, we give upper bounds on the sensitivity of $\chi$ to string operations in terms of measure $\delta$.
@@
$\chi(w^{\prime})-\chi(w)=O\left(\delta\max\left(1,\log(n/(\delta\log\delta))\right)\log\delta\right)\text{ and }$
$\chi(w^{\prime})\hskip 3.0pt/\hskip 4.0pt\chi(w)=O\left(\max\left(1,\log(n/(\delta\log\delta))\right)\log\delta\right).$
@@
It holds $\delta\leq\chi\leq 2\overline{r}$ [4] and $r=O(\delta\max(1,\log(n/(\delta\log\delta)))\log\delta)$ [17]. Since the multiplicative sensitivity of $\delta$ to any of the string operations is $O(1)$ [1], for any $w\in\Sigma^{*}$ it holds $\overline{r}(w)=r(w^{R})=O(\delta\max(1,\log(n/(\delta\log\delta)))\log\delta)$. The thesis follows by considering the worst case, that is, $\chi(w)=\Theta(\delta)$ and $\chi(w^{\prime})=\Theta(\delta\max(1,\log(n/(\delta\log\delta)))\log\delta)$.∎
@@ 5.1 Proving $\chi=O(r)$
Let $x_{i}$ denotes the $i$th rotation of $w\text{\tt\textdollar}$ in lexicographic order, for each $i\in[1\mathinner{.\,.}|w|+1]$, and let $u_{i}$ be the longest common prefix between the rotations $x_{i},x_{i+1}$, for each $i\in[1\mathinner{.\,.}|w|]$ (note this implies that $u_{i}$ is right-maximal). We further define $s:[1\mathinner{.\,.}n+1]\rightarrow[0\mathinner{.\,.}n]$ as $s(i)=j$ if $x_{i}=w[j+1\mathinner{.\,.}|w|]\text{\tt\textdollar}w[1\mathinner{.\,.}j]$, that is, the number of cyclic shifts to the right required to transform $x_{i}$ into $w\text{\tt\textdollar}$. [^2] The function $s$ mimics the well-known Suffix Array [23], here omitted for simplicity of exposition. As the symbol \$ occurs only once in $w\text{\tt\textdollar}$, the function $s$ is bijective.
@@
that is, the set of positions where the occurrences of the right-extensions $u_{i}c_{1}$ and $u_{i}c_{2}$ end in $w\text{\tt\textdollar}$, where $u_{i}c_{1}$ and $u_{i}c_{2}$ are the prefixes of $x_{i}$ and $x_{i+1}$, respectively, for some $c_{1},c_{2}\in\Sigma$ such that $c_{1}<c_{2}$. It follows by construction that the set $S$ is a suffixient set of $w\text{\tt\textdollar}$.
@@ 5.2 A family with $\chi=o(v)$ (and thus $o(r)$)
An infinite string w is episturmian if it has at most one right-maximal substring of each length and its set of substrings is closed under reversal, that is, $\mathcal{F}_{\textbf{w}}=\mathcal{F}_{\textbf{w}}^{R}$. It is standard episturmian (or epistandard) if, in addition, all the right-maximal substrings of w are of the form $\textbf{w}[1\mathinner{.\,.}i]^{R}$ with $i\geq 0$, that is, they are the reverse of some prefix of w.
@@
Let $\textbf{w}\in\Sigma^{\omega}$ be an episturmian word with $|\Sigma|=\sigma\geq 2$. Then, $\mathtt{sre}(\textbf{w}[i\mathinner{.\,.}j])\leq\sigma$ for any $i,j\geq 0$.
@@
Let w be an epistandard word. The right-extensions $x_{1},x_{2},\dots$ ending with $a\in\Sigma$ form a suffix-chain where each $x_{i}$ is a suffix of $x_{i+1}$. This is because the right-maximal substrings are the reverses of the prefixes of the word, so they are all suffixes of the longest one. If we add the same letter to a subset of them to obtain right-extensions, they still form a suffix-chain. There is one of those suffix-chains for each character $a\in\Sigma$.
@@
When considering substrings of w, the super-maximal right-extension in $\textbf{w}[i\mathinner{.\,.}j]$ ending with $a\in\Sigma$ is the longest right-extension of w ending with $a$ that remains a right-extension in $\textbf{w}[i\mathinner{.\,.}j]$. It follows that for any substring $\textbf{w}[i\mathinner{.\,.}j]$ of any episturmian word w, it holds $\mathtt{sre}(\textbf{w}[i\mathinner{.\,.}j])\leq\sigma$. ∎
@@ 6 Online Computation of Smallest Suffixient Sets
As an application of Lemmas 2 and 3, we show that Ukkonen’s [33] and Weiner’s [34] linear-time online constructions of suffix trees can be easily modified to compute smallest suffixient sets within the same space and time complexity.
The suffix tree of a text $T$ is a tree of size $O(|T|)$ where edges are labeled by nonempty substrings of $T$ (the label $T[i\mathinner{.\,.}j]$ is indicated by the pair $(i,j)$). Every internal node has at least two children, and the labels of edges to any two children must differ in their first symbol. If node $x$ has a child node $y$ by an edge labeled $(i,j)$, we say that $y$ is a child by label $T[i]$. Each node $x$ is said to be the locus of the string obtained by concatenating the labels of the edges that lead from the root to $x$. The string depth of a node is the length of its locus. The tree leaves are the loci of suffixes $T[i\mathinner{.\,.}]$, whereas internal nodes are loci of strings that occur at least twice in $T$. Edges to leaves have labels of the form $(i,\infty)$, which means the second component is always the last position processed of $T$. Suffix tree nodes also have so-called suffix links, which lead from a node that is the locus of a string $cu$, for $c\in\Sigma$, to the locus of string $u$ (which always exists: if $cu$ occurs more than once, so does $u$). Finally, we extend the concept of suffix tree nodes to implicit nodes, which are virtual nodes with one child, assumed to exist along edges: if $y$ is the child of $x$ by an edge labeled $(i,j)$ and $x$ is the locus of string $u$, then $(x,(i,p))$, for $i\leq p<j$, is an implicit node that is the locus of $u\cdot T[i\mathinner{.\,.}p]$. Knowing the loci of which string they are, suffix links are also defined from implicit nodes (those can lead to explicit or to implicit nodes).
6.1 Ukkonen’s left-to-right construction
Ukkonen’s algorithm [33] builds the suffix tree of $T$ such that, after having processed any prefix $w$ of $T$, it has built the suffix tree of $w$. To prevent special cases, it assumes that the root is in fact a child of a special node $\bot$, with edges to the root labeled $c$ for every $c\in\Sigma$; the suffix link of the root points to $\bot$. We do not aim at a full explanation of the algorithm, but just highlight some of its relevant properties. Let the prefix $w$ of $T$ be followed by $c\in\Sigma$. The algorithm maintains pointers to two (possibly implicit) nodes of the suffix tree of $w$: